# Totally elliptic surface group representations

### v1

@@ 2.3.3. The case $n\geq 5$
Let now $\Sigma$ denote a sphere with $n\geq 5$ punctures and $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ be a reduced totally elliptic representation. We’ll see that the proof of Theorem A in this case can be brought back to the case of 4-punctured spheres (Section 2.3.2). As before, we start with the case of representations with a non-peripheral simple closed curve in their kernel.
@@
We choose a geometric generating family of $\pi_{1}\Sigma$ consisting of generators $c_{1},\ldots,c_{n}$ satisfying $c_{1}\cdots c_{n}=1$. They can be chosen in a way that the simple closed curve in the kernel of $\rho$ is represented by the fundamental group element $c_{1}\cdots c_{i+1}$ for some $i\in\{2,\ldots,n-3\}$. This implies that $\rho(c_{i+1})=\rho(c_{1}\cdots c_{i})^{-1}$ and $\rho(c_{i+2})=\rho(c_{i+3}\cdots c_{n})^{-1}$. Arguing as in the proof of Lemma 2.13, we know that we can conjugate $\rho$ so that $\rho(c_{i+1})$ and $\rho(c_{i+2})$ both belong to $\operatorname{PSO}(2)$.
@@
Now, note that the fundamental group element $c_{i+1}c_{j}(c_{1}\cdots c_{i})c_{j}^{-1}$ represents a simple closed curve for every $j\geq i+2$ and similarly for $c_{i+2}^{-1}c_{j}^{-1}(c_{i+3}\cdots c_{n})^{-1}c_{j}$ when $j\leq i+1$. The first one is the image of $c_{1}\cdots c_{i+1}$ by the Dehn twist along $c_{i+1}c_{j}$ and the second one is the image of $(c_{i+2}\cdots c_{n})^{-1}$ by the Dehn twist along $c_{j}c_{i+2}$. Since $\rho$ is totally elliptic, we conclude that $\rho(c_{i+1}c_{j}(c_{1}\cdots c_{i})c_{j}^{-1})=[\rho(c_{i+1}),\rho(c_{j})]$ and $\rho(c_{i+2}^{-1}c_{j}^{-1}(c_{i+3}\cdots c_{n})^{-1}c_{j})=[\rho(c_{i+2})^{-1},\rho(c_{j})^{-1}]$ are both elliptic, hence trivial by Lemma 2.1. Because we’re assuming $\rho$ to be reduced, it holds that $\rho(c_{i+1})\neq 1$ and $\rho(c_{i+2})\neq 1$. We conclude that $\rho(c_{j})$ belongs to $\operatorname{PSO}(2)$ for every $j$ again by Lemma 2.1. This proves that $\rho$ is orthogonal. ∎
@@
From now on, let’s assume that $\rho$ is a reduced totally elliptic representation with no simple closed curves in its kernel ($\rho$ is “regularly totally elliptic” in the terminology introduced after Definition 2.10). To such a representation corresponds a chain of triangles in the hyperbolic plane $\operatorname{\mathbb{H}}$. It’s constructed from a chained pants decomposition $\mathcal{B}$ of $\Sigma$, by which we mean a pants decomposition of $\Sigma$ in which each pair of pants contains at least one puncture. We then choose a geometric presentation of $\pi_{1}\Sigma$ with generators $c_{1},\ldots,c_{n}$ satisfying $c_{1}\cdots c_{n}=1$, such that the $n-3$ pants curves of $\mathcal{B}$ are represented by the fundamental group elements $b_{i}=(c_{1}\cdots c_{i+1})^{-1}$ for $i=1,\ldots,n-3$. This is always possible as explained in [Mar24, Appendix B]. Since $\rho$ is reduced and has no simple closed curves in its kernel, all the elliptic elements $\rho(c_{1}),\ldots,\rho(c_{n})$ and $\rho(b_{1}),\ldots,\rho(b_{n-3})$ are regular, and have a unique fixed point in $\operatorname{\mathbb{H}}$ which we denote by $C_{1},\ldots,C_{n}$ and $B_{1},\ldots,B_{n-3}$. The triangle chain of $\rho$ consists of the $n-2$ triangles
@@
If there exists a chained pants decomposition $\mathcal{B}$ such that all the triangles in the $\mathcal{B}$-triangle chain of $\rho$ are non-degenerate and have the same orientation, then $\rho$ is a DT representation. Conversely, if $\rho$ is a DT representation, then for every chained pants decomposition $\mathcal{B}$, all non-degenerate triangles in the $\mathcal{B}$-triangle chain of $\rho$ have the same orientation.
If the $\mathcal{B}$-triangle chain of $\rho$ only contains degenerate triangles, then all $\rho(c_{i})$ fix the same point. This means that $\rho$ is an orthogonal representation. Instead, if at least one triangle is non-degenerate, then the elliptic elements $\rho(c_{i})$ are not all rotations around the same point in $\operatorname{\mathbb{H}}$ and $\rho$ is not orthogonal. In the latter case, we can actually replace $\mathcal{B}$ by a new chained pants decomposition of $\Sigma$ and find a compatible geometric presentation of $\pi_{1}\Sigma$ such that the new triangle chain of $\rho$ only consists of non-degenerate triangles. This new pants decomposition can be constructed using the same algorithm as in [FM23, Proposition 2] and the compatible geometric presentation of $\pi_{1}\Sigma$ is obtained by the same reasoning as in [FM23, Proposition 3]. From now on, we’ll therefore assume that all the triangles in the $\mathcal{B}$-triangle chain of $\rho$ are non-degenerate.
We’ll denote the rotation angles of the images of the pants curve $b_{1},\ldots,b_{n-3}$ in $\mathcal{B}$ under $\rho$ by $\beta_{1},\ldots,\beta_{n-3}\in(0,2\pi)$. Recall that $\rho(b_{1}),\ldots,\rho(b_{n-3})$ are all regular elliptic elements because we’re assuming that the kernel of $\rho$ doesn’t contain any simple closed curve.
@@
If $\rho$ is a reduced totally elliptic representation with no simple closed curves in its kernel, then all the triangles in the $\mathcal{B}$-triangle chain of $\rho$ have the same orientation. Hence, $\rho$ is a DT representation by Proposition 2.16.
@@
$\operatorname{Rep}_{\alpha^{(i)}}(\Sigma^{(i)},\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$
@@
The standard pants decomposition of $\Sigma^{(i)}$ associated to the geometric system of generators $((b_{i-1})^{-1},c_{i+1},c_{i+2},b_{i+1})$ of $\pi_{1}\Sigma^{(i)}$ consists of one curve: $b_{i}$. We’ll write it $\mathcal{B}_{i}$. The $\mathcal{B}_{i}$-triangle chain of the restriction of $\rho$ to $\Sigma^{(i)}$ is made of the two triangles that share the vertex $B_{i}$ in the $\mathcal{B}$-triangle chain of $\rho$. These two triangles are non-degenerate because we’re assuming that every triangle in the $\mathcal{B}$-triangle chain of $\rho$ is non-degenerate. In particular, this means that none of the restrictions of $\rho$ to any of the spheres $\Sigma^{(i)}$ is a representation into $\operatorname{PSO}(2)$. So, by the discussion of Section 2.3.2 about 4-punctured spheres, they’re all DT representations because they are totally elliptic. This means that the two triangles in the $\mathcal{B}_{i}$-triangle chain of $\rho$ have the same orientation by Proposition 2.16. Now, because the $\mathcal{B}$-triangle chain of $\rho$ is made of the $\mathcal{B}_{i}$-triangle chain of each restriction, with an overlap of one non-degenerate triangle for consecutive indices, we conclude that all the triangles in the $\mathcal{B}$-triangle chain of $\rho$ have the same orientation. ∎

### v2

@@ 2.3.3. The case $n\geq 5$
Let now $\Sigma$ denote a sphere with $n\geq 5$ punctures and $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ be a reduced totally elliptic representation. As before, we start with the case of representations with a non-peripheral simple closed curve in their kernel.
@@
We choose a geometric generating family of $\pi_{1}\Sigma$ consisting of generators $c_{1},\ldots,c_{n}$ satisfying $c_{1}\cdots c_{n}=1$, similarly as on Figure 1. They can be chosen in a way that the simple closed curve in the kernel of $\rho$ is represented by the fundamental group element $c_{1}\cdots c_{i+1}$ for some $i\in\{1,\ldots,n-3\}$. This implies that $\rho(c_{i+1})=\rho(c_{1}\cdots c_{i})^{-1}$ and $\rho(c_{i+2})=\rho(c_{i+3}\cdots c_{n})^{-1}$. We will argue as in the proof of Lemma 2.13. First, note that the fundamental group element $(c_{1}\cdots c_{i})c_{i+2}^{-1}c_{i+1}c_{i+2}$ represents a simple closed curve on $\Sigma$ that is the image of the simple closed curve represented by $c_{1}\cdots c_{i+1}$ under the Dehn twist along $c_{i+1}c_{i+2}$. Since we are assuming $\rho$ to be totally elliptic, it implies that $\rho\big{(}(c_{1}\cdots c_{i})c_{i+2}^{-1}c_{i+1}c_{i+2}\big{)}=\big{[}\rho(c_{i+1}^{-1}),\rho(c_{i+2}^{-1})\big{]}$ is elliptic. We deduce that $\big{[}\rho(c_{i+1}),\rho(c_{i+2})\big{]}=1$ from Lemma 2.1. So, up to conjugating $\rho$, we may assume that $\rho(c_{i+1})$ and $\rho(c_{i+2})$ both belong to $\operatorname{PSO}(2)$.
@@
Now, observe further that the fundamental group element $(c_{1}\cdots c_{i})c_{j}^{-1}c_{i+1}c_{j}$ represents a simple closed curve for every $j\geq i+2$ which is the image of the simple closed curve represented by $c_{1}\cdots c_{i+1}$ under the Dehn twist along $c_{i+1}c_{j}$. Similarly, for every $j\leq i+1$, $(c_{i+3}\cdots c_{n})^{-1}c_{j}c_{i+2}^{-1}c_{j}^{-1}$ represents the simple closed curve obtained by applying the Dehn twist along $c_{j}c_{i+2}$ to the simple closed curve represented by $(c_{i+2}\cdots c_{n})^{-1}$. Since $\rho$ is totally elliptic, we conclude that $\rho\big{(}(c_{1}\cdots c_{i})c_{j}^{-1}c_{i+1}c_{j}\big{)}=\big{[}\rho(c_{i+1})^{-1},\rho(c_{j})^{-1}\big{]}$ and $\rho\big{(}(c_{i+3}\cdots c_{n})^{-1}c_{j}c_{i+2}^{-1}c_{j}^{-1}\big{)}=\big{[}\rho(c_{i+2}),\rho(c_{j})\big{]}$ are both elliptic, hence trivial by Lemma 2.1. Because we are assuming $\rho$ to be reduced, it holds that $\rho(c_{i+1})\neq 1$ and $\rho(c_{i+2})\neq 1$. We conclude that $\rho(c_{j})$ belongs to $\operatorname{PSO}(2)$ for every index $j$. This proves that $\rho$ is orthogonal. ∎
@@
From now on, we will assume that $\rho$ is a reduced totally elliptic representation with no simple closed curves in its kernel. In the terminology introduced after Definition 2.10, $\rho$ is regularly totally elliptic. To such a representation corresponds a chain of triangles in the hyperbolic plane $\operatorname{\mathbb{H}}$. It is constructed from a chained pants decomposition $\mathcal{B}$ of $\Sigma$, by which we mean a pants decomposition of $\Sigma$ in which each pair of pants contains at least one puncture. We then choose a geometric presentation of $\pi_{1}\Sigma$ with generators $c_{1},\ldots,c_{n}$ satisfying $c_{1}\cdots c_{n}=1$, such that the $n-3$ pants curves of $\mathcal{B}$ are represented by the fundamental group elements $b_{i}=(c_{1}\cdots c_{i+1})^{-1}$ for $i=1,\ldots,n-3$. This is always possible as explained in [Mar24, Appendix B]. Since $\rho$ is reduced and has no simple closed curves in its kernel, all the elliptic elements $\rho(c_{1}),\ldots,\rho(c_{n})$ and $\rho(b_{1}),\ldots,\rho(b_{n-3})$ are regular, and have a unique fixed point in $\operatorname{\mathbb{H}}$ which we denote by $C_{1},\ldots,C_{n}$ and $B_{1},\ldots,B_{n-3}$. The triangle chain of $\rho$ consists of the $n-2$ triangles
@@
Let $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ be a reduced regularly totally elliptic representation. If there exists a chained pants decomposition $\mathcal{B}$ of $\Sigma$ such that the $\mathcal{B}$-triangle chain of $\rho$ contains at least one non-degenerate triangle and all non-degenerate triangles have the same orientation, then $\rho$ is a DT representation. Conversely, if $\rho$ is a DT representation, then for every chained pants decomposition $\mathcal{B}$ of $\Sigma$, all non-degenerate triangles in the $\mathcal{B}$-triangle chain of $\rho$ have the same orientation.
If the $\mathcal{B}$-triangle chain of $\rho$ only contains degenerate triangles, then all $\rho(c_{i})$ fix the same point. This means that $\rho$ is an orthogonal representation. Instead, if at least one triangle is non-degenerate, then the elliptic elements $\rho(c_{i})$ are not all rotations around the same point and $\rho$ is not orthogonal. In the latter case, we can actually replace $\mathcal{B}$ by a new chained pants decomposition of $\Sigma$ and find a compatible geometric presentation of $\pi_{1}\Sigma$ such that the new triangle chain of $\rho$ only consists of non-degenerate triangles. This new pants decomposition can be constructed using the same algorithm as in [FM23, Proposition 2] and the compatible geometric presentation of $\pi_{1}\Sigma$ is obtained by the same reasoning as in [FM23, Proposition 3]. From now on, we will therefore assume that all the triangles in the $\mathcal{B}$-triangle chain of $\rho$ are non-degenerate.
We denote the rotation angles of the images of the pants curve $b_{1},\ldots,b_{n-3}$ in $\mathcal{B}$ under $\rho$ by $\beta_{1},\ldots,\beta_{n-3}\in(0,2\pi)$. Recall that $\rho(b_{1}),\ldots,\rho(b_{n-3})$ are all regular elliptic because we are assuming $\rho$ to be regularly totally elliptic.
@@
Under our assumptions that $\rho$ is a reduced regularly totally elliptic representation and that its $\mathcal{B}$-triangle chain is made of non-degenerate triangles, then all triangles must have the same orientation. Hence, $\rho$ is a DT representation by Proposition 2.17.
@@
$\operatorname{Rep}_{\alpha^{(i)}}(\Sigma^{(i)},\operatorname{PSL}_{2}\operatorname{\mathbb{R}}),$
@@
The standard pants decomposition of $\Sigma^{(i)}$ associated to the geometric system of generators $((b_{i-1})^{-1},c_{i+1},c_{i+2},b_{i+1})$ of $\pi_{1}\Sigma^{(i)}$ consists of one simple closed curve represented by $b_{i}$. We will denote this pants decomposition by $\mathcal{B}_{i}$. The $\mathcal{B}_{i}$-triangle chain of the restriction of $\rho$ to $\Sigma^{(i)}$ is made of the two triangles that share the vertex $B_{i}$ in the $\mathcal{B}$-triangle chain of $\rho$. These two triangles are non-degenerate by assumption. In particular, this means that none of the restrictions of $\rho$ to any of the spheres $\Sigma^{(i)}$ is orthogonal. So, by the discussion of Section 2.3.2 about 4-punctured spheres, all the restrictions are DT representations because they are totally elliptic. This means that the two triangles in the $\mathcal{B}_{i}$-triangle chain of $\rho$ have the same orientation by Proposition 2.17. Now, because the $\mathcal{B}$-triangle chain of $\rho$ is made of the $\mathcal{B}_{i}$-triangle chain of each restriction, with an overlap of one non-degenerate triangle for consecutive indices, we conclude that all the triangles in the $\mathcal{B}$-triangle chain of $\rho$ have the same orientation. ∎
@@
3. Totally elliptic representations in $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$
3.1. Overview
The goal of this section is to classify totally elliptic surface group representations into $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$—the complexification of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$—by proving Theorems C and D. We will first prove that irreducible totally elliptic representations into $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ are either unitary or DT representations (Section 3.2), and then consider the case of reducible totally elliptic representations (Section 3.3).
3.2. Irreducible representations
We start with the case of irreducible surface group representations into $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ which are, by definition, representations whose image is not conjugate to a subgroup of $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ made of upper triangular elements.
Proposition 3.1.
Let $\rho\colon\pi_{1}\Sigma\to\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$ be a reduced totally elliptic surface group representation. If $\rho$ is irreducible, then either $\rho$ is unitary, or it is conjugate to the composition of a DT representation with the inclusion $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}\subset\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$.
Proof.
For every $\gamma\in\pi_{1}\Sigma$, choose a representative $A_{\gamma}\in\mathrm{SL}_{2}\operatorname{\mathbb{C}}$ of $\rho(\gamma)$ and write $t_{\gamma}$ for the trace of $A_{\gamma}$. Since $\rho$ is totally elliptic, for every $\sigma\in\pi_{1}\Sigma$ that represents a simple closed curve on $\Sigma$, the number $t_{\sigma}$ is real. Using trace relations for $2\times 2$ matrices, it is actually possible to write every trace $t_{\gamma}$ as a polynomial expression in the traces $t_{\sigma}$, where $\sigma$ ranges over fundamental group elements representing simple closed curves. This was initially observed by Goldman–Xia in their work on ergodic dynamics on character varieties [GX11, Section 2] as a consequence of classical results by Procesi [Pro76]. In particular, this means that the traces $t_{\gamma}$ are real numbers for all $\gamma\in\pi_{1}\Sigma$. Since $\rho$ is irreducible, it implies that the image of $\rho$ is conjugate into a real form of $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ (see for instance [MR03, Corollary 3.2.5]). The real forms of $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ being $\operatorname{PSU}(2)$ and $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, we conclude that $\rho$ is either unitary or conjugate to a representation into $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. In the latter case, since $\rho$ is reduced, Theorem A applies and shows that $\rho$ is conjugate to the composition of a DT representation with the inclusion $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}\subset\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$. ∎
3.3. Reducible representations
A reducible surface group representation $\rho\colon\pi_{1}\Sigma\to\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$ is conjugate to a representation whose image is contained in the subgroup of upper triangular elements of $\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$. So, up to conjugating a reducible representation $\rho$, we may decompose it into a linear part $\lambda\colon\pi_{1}\Sigma\to\operatorname{\mathbb{C}}^{\times}$ and a function $z\colon\pi_{1}\Sigma\to\operatorname{\mathbb{C}}$ by writing
$\rho(\gamma)=\pm\begin{pmatrix}\lambda(\gamma)&\lambda(\gamma)^{-1}z(\gamma)\\ 0&\lambda(\gamma)^{-1}\end{pmatrix}.$
The function $z$ turns out to be a cocycle in the sense that $z(\gamma_{1}\gamma_{2})=z(\gamma_{1})+\lambda(\gamma_{1})^{2}z(\gamma_{2})$ for every $\gamma_{1},\gamma_{2}\in\pi_{1}\Sigma$. If $\rho$ is totally elliptic, then $|\lambda(\gamma)|=1$ for every $\gamma\in\pi_{1}\Sigma$ because any system of geometric generators of $\pi_{1}\Sigma$ is made of simple closed curves (as the one illustrated on Figure 1).
Proposition 3.2.
Consider a surface $\Sigma$ of genus $g\geq 1$ and with $n\geq 0$ punctures, and let $\rho\colon\pi_{1}\Sigma\to\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$ be a reduced totally elliptic representation. If $\rho$ is reducible, then the image of $\rho$ is conjugate to a subgroup made of diagonal elements in $\operatorname{PSU}(2)$.
Proof.
Up to conjugating $\rho$, we may assume it is upper triangular. A simple computation shows that the commutator of two upper triangular $2\times 2$ matrices is elliptic if and only if it is trivial. More precisely, it holds that
(3.1) $\left[\begin{pmatrix}\lambda(\gamma_{1})&\lambda(\gamma_{1})^{-1}z(\gamma_{1})\\ 0&\lambda(\gamma_{1})^{-1}\end{pmatrix},\begin{pmatrix}\lambda(\gamma_{2})&\lambda(\gamma_{2})^{-1}z(\gamma_{2})\\ 0&\lambda(\gamma_{2})^{-1}\end{pmatrix}\right]=\begin{pmatrix}1&z(\gamma_{1}\gamma_{2})-z(\gamma_{2}\gamma_{1})\\ 0&1\end{pmatrix}.$
This statement is an analogue of Lemma 2.1. It means that we can apply the arguments of Section 2.2 again to conclude that $\rho$ is an abelian representation. Now, if $\rho$ is abelian and upper triangular, then it is conjugate to the representation $\rho^{\prime}\colon\pi_{1}\Sigma\to\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$ given by
$\rho^{\prime}(\gamma)=\pm\begin{pmatrix}\lambda(\gamma)&0\\ 0&\lambda(\gamma)^{-1}\end{pmatrix}.$
Since $\rho$ is totally elliptic, it holds that $|\lambda(\gamma)|=1$ for every $\gamma\in\pi_{1}\Sigma$ and thus $\rho^{\prime}$ is unitary. ∎
Propositions 3.1 and 3.2 together prove Theorem C. We now proceed to the proof of Theorem D.
The reason why Theorem C does not apply to punctured spheres is because there exist reduced totally elliptic representations $\pi_{1}\Sigma\to\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$ that are reducible but not unitary when $\Sigma$ is a sphere with $n\geq 3$ punctures. Before describing such examples, we start with the following result about simple closed curves on $\Sigma$.
Lemma 3.3.
Let $\Sigma$ be a sphere with $n\geq 1$ punctures and $c_{1},\ldots,c_{n}$ be a system of geometric generators of $\pi_{1}\Sigma$ (like those depicted on Figure 1). If $\gamma\in\pi_{1}\Sigma$ represents a simple closed curve on $\Sigma$, then there exist $k\geq 1$ and $1\leq i_{1}<\cdots<i_{k}\leq n$ such that either $\gamma$ or $\gamma^{-1}$ is equal to $c_{i_{1}}\cdots c_{i_{k}}$ in the abelianization of $\pi_{1}\Sigma$.
Proof.
The group $\mathrm{Aut}^{\ast}(\pi_{1}\Sigma)$ consists of all automorphisms of $\pi_{1}\Sigma$ that map each generator $c_{i}$ to a conjugate of itself. In particular, two fundamental group elements that are images of each other by an automorphism in $\mathrm{Aut}^{\ast}(\pi_{1}\Sigma)$ coincide in the abelianization of $\pi_{1}\Sigma$. The image of $\mathrm{Aut}^{\ast}(\pi_{1}\Sigma)$ inside the group of outer automorphisms of $\pi_{1}\Sigma$ is isomorphic to the pure mapping class group of $\Sigma$, which means that $\mathrm{Aut}^{\ast}(\pi_{1}\Sigma)$ preserves the subset of $\pi_{1}\Sigma$ made of all fundamental group elements representing simple closed curves on $\Sigma$. Moreover, if $\gamma\in\pi_{1}\Sigma$ is a fundamental group element representing a simple closed curve, then the union of the $\mathrm{Aut}^{\ast}(\pi_{1}\Sigma)$-orbits of $\gamma$ and $\gamma^{-1}$ contains exactly one product of the form $c_{i_{1}}\cdots c_{i_{k}}$ for $1\leq i_{1}<\cdots<i_{k}\leq n$ and $11\leq k\leq n-$ 1. This concludes the proof. ∎
Example 3.4.
Using the same notation as in Lemma 3.3, we consider a representation $\rho\colon\pi_{1}\Sigma\to\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$ of a sphere $\Sigma$ with $n\geq 3$ punctures whose image is upper triangular. As above, we denote the image of $\gamma\in\pi_{1}\Sigma$ by
$\rho(\gamma)=\pm\begin{pmatrix}\lambda(\gamma)&\lambda(\gamma)^{-1}z(\gamma)\\ 0&\lambda(\gamma)^{-1}\end{pmatrix}.$
We claim that if the linear part $\lambda\colon\pi_{1}\Sigma\to\operatorname{\mathbb{C}}^{\times}$ is chosen such that
(3.2) $|\lambda(c_{i_{1}}\cdots c_{i_{k}})|=1\quad\text{and}\quad\lambda(c_{i_{1}}\cdots c_{i_{k}})\neq\pm 1$
for every $1\leq i_{1}<\cdots<i_{k}\leq c_{n}$ with $11\leq k\leq n-$ 1, then $\rho$ is a reduced totally elliptic representation. We shall now prove this claim. First, observe that $\rho$ is reduced because $\lambda(c_{i})\neq\pm 1$ for every $i=1,\ldots,n$. Now, if $\gamma\in\pi_{1}\Sigma$ represents a simple closed curve on $\Sigma$, then $\lambda(\gamma)=\lambda(c_{i_{1}}\cdots c_{i_{k}})^{\pm 1}$ for some $1\leq i_{1}<\cdots<i_{k}\leq n$ with $11\leq k\leq n-$ 1 by Lemma 3.3. Since by assumption $|\lambda(c_{i_{1}}\cdots c_{i_{k}})|=1$ and $\lambda(c_{i_{1}}\cdots c_{i_{k}})\neq\pm 1$, we conclude that $\rho$ is totally elliptic.
Let us now prove that for every $\lambda$ satisfying (3.2), we can always find an assorted cocycle $z\colon\pi_{1}\Sigma\to\operatorname{\mathbb{C}}$ such that $\rho$ is not unitary. Since $\rho$ is upper triangular, the image of $\rho$ has a global fixed point for the action of $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ on $\mathbb{CP}^{1}$, seen here as the boundary of the 3-dimensional hyperbolic space. This means that $\rho$ is unitary if and only if its image fixes every point on a geodesic line in the 3-dimensional hyperbolic space, which would imply that $\rho$ is abelian. So, if we choose the cocycle $z$ such that there exist $i\neq j$ with $z(c_{i}c_{j})\neq z(c_{j}c_{i})$, then $\rho$ is not abelian because of (3.1), hence not unitary. Such cocycles $z$ always exist as soon as $n\geq 3$.
In conclusion, by picking $\lambda$ and $z$ as we just described, we obtain a reduced totally elliptic representation $\rho\colon\pi_{1}\Sigma\to\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$ that is reducible but not unitary. This finishes the proof of Theorem D.